4. Assume that the mean life of a particular brand of car battery is normally distributed with a mean of 28 months and a standard variation of 4 months.


(a) for a randomly selected battery of this make, what is the probability that it will last between 30 and 34 months?.

(b) what is the probability that a randomly selected battery of this make will fail within years of date of purchase?.


(c) after what time period will 60% of all batteries of this makes fail?.


1
Expert's answer
2021-01-28T05:26:24-0500

We denote the respective random variable (life of battery) by "X". We remind that for the normal distribution with parameters "\mu=28" and "\sigma=4" the probability density function is "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^2}"

(a) "P(30\leq X\leq34)=\int_{30}^{34}p(x)dx=\frac{1}{\sigma\sqrt{2\pi}}\int_{30}^{34}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^2}dx\approx0.2417"

We computed the integral using the following code in Anaconda(https://www.anaconda.com/):


from scipy import integrate

import numpy as np

import math

func = lambda x:(1/(4*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-28)/4)*((x-28)/4))

e = integrate.quad(func, 30, 34)

print(e)


(b) We assume that the time is 1 year. We get: "P(-\infty\leq X\leq12)=\int_{-\infty}^{12}p(x)dx=\frac{1}{\sigma\sqrt{2\pi}}\int_{-\infty}^{12}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^2}dx\approx0.00003167"

(c) The aim is to find such "\alpha" that: "P(-\infty\leq X\leq\alpha)=\frac{1}{\sigma\sqrt{2\pi}}\int_{-\infty}^{\alpha}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^2}dx=0.4" It means that the battery will fail with the probability 60%. We get that "\alpha\approx26.99". We received "\alpha" by substituting different values in the code (Anaconda):


from scipy import integrate

import numpy as np

import math

func = lambda x:(1/(4*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-28)/4)*((x-28)/4))

e = integrate.quad(func,  -np.inf, 26.99)

print(e)